OpenStudy (anonymous):

Can someone help me please? y-1 3/10 = 4 5/6

OpenStudy (anonymous):

so the is Y-1.3= 7.5 ?

OpenStudy (xishem):

I'm pretty sure it's:\[y-1\frac{3}{10}=4\frac{5}{6}\]

OpenStudy (anonymous):

just add 1 3/10 to the other side of the equation to solve for y.

OpenStudy (anonymous):

so you get \[y =6\frac{ 2 }{ 15 }\]

OpenStudy (anonymous):